OpenStudy (buttermequeasy):

Choose the slope-intercept equation of the line that passes through the point (6, 2) and is perpendicular to y = 3x + 6. y = one thirdx y = -one thirdx + 4 y = 3x + 20 y = -3x - 16

OpenStudy (buttermequeasy):

help!

OpenStudy (anonymous):

did you get slope of line ?

OpenStudy (callisto):

Since the line is perpendicular to y = 3x + 6 slope of line y = 3x + 6 is 3 (Luis_Rivera has written that for you) therefore slope of line x 3 = -1 slope of line = -1/3 Equation is (y-y1) = m (x-x1) Pur (x1, y1) = (6,2) and m = -1/3 into the equation y - 2 = (-1/3) (x -6) y = (-x/3) + 2 +2 y = -x/3 +4 So choose the suitable answer!!

OpenStudy (anonymous):

m=-1/3 y-y1=m(x-x1) y-2=-1/3(x-6) y-2=-1/3x+2 y=-1/3x+4